What is the value of Kc for the equilibrium below at 300 K?

$$

1

2

​A2

​(g)+1

2

​B2

​(g)AB(g)



1
Expert's answer
2021-11-23T13:16:01-0500

A2(g)+B2(g) ↔ 2AB(g)    kc = 3.65

1/2A2(g)+1/2B2(g↔AB(g)    kc1 = ?

The values in the equilibrium equation are halved so the new equilibrium constant is the square root of the previous value;

kc1 = (kc)1/2 = (3.65)1/2 = 1.91